OpenStudy (loser66):

Please, help Determine X_n converges or diverges \(X_n=\dfrac{(-1)^n*n}{n+1}\)

OpenStudy (loser66):

I have solution here. What I don't know is how to link \(X_n\) and \(X_{n+2} \) to complete the proof.

OpenStudy (loser66):

@ganeshie8

ganeshie8 (ganeshie8):

Clearly the two subsequences are convergent

ganeshie8 (ganeshie8):

since they are converging to different values, the overall sequence wont converge

ganeshie8 (ganeshie8):

as n goes to infinity, the even terms are converging to 1 and the odd terms are converging to -1

OpenStudy (loser66):

oh, use convergent subsequences.

ganeshie8 (ganeshie8):

so you will see the terms as : ...., 0.98, -0.98, 0.99, -0.99,...

ganeshie8 (ganeshie8):

so the soverall sequence is dancing between -1 and 1forever

OpenStudy (loser66):

Can I use limit theorem to show it diverges? For this proof, they use subsequence, but the problem asks me to use limit theorem, which is : if Xn = Yn *Zn, and BOTH Yn, Zn converge, then Xn converges I have Yn =(-1)^n diverges Zn = \(\dfrac{n}{n+1}\) converges to 1 ( I can prove it) But I don't know whether I can apply that theorem to state that Xn diverges

ganeshie8 (ganeshie8):

yeah its one way i think : if Yn and Zn converge, then Yn*Zn converges

ganeshie8 (ganeshie8):

if Yn diverges and Zn converges, then we `don't know` if Yn*Zn diverges or not.

OpenStudy (loser66):

Got you. Thank you very much. I make a lot of questions to make sure that I understand the concept completely. :) Thanks for being patient to me. :)

ganeshie8 (ganeshie8):

thats good only :) here is an example : Yn = n diverges Zn = 1/n^2 converges but the sequence Yn*Zn = 1/n converges.

OpenStudy (loser66):

:)